The following data are available relating to the performance of Seminole Fund and the market portfolio: Market Seminole PortfMultiple Choice o 4.0% o 20.0% o 2.86% o 0.8% o 40.0%
Answer #1

M2 measure is also called risk adjusted performance and is an improvement over sharpe ratio.

Following is the formula of M2 =(Rp - Rf) * \sigma m / \sigma p - (Rm - Rf)

Whrere, Rp =Return on Seminole Fund (14%), Rf = Risk free rate of return(6%), Rm= Return on Market Portfolio(18%),

\sigmam = Standard Deviation of return on market portfolio (22%), \sigma p= Standard Deviation of return on fund (30%)

So, by putting values in the above formula we get,

M2= (18 - 6) * 22 / 30 - (14 - 6)

= (12 * 0.73 ) - 8

= 8.8 - 8

= 0.8 i.e 0.8%

Therefore 0.8% is the correct answer.

Note: The other options given is incorrect because the given option has been arrived on the basis of formula.